# AED: Automatic Discovery of Effective and Diverse Vulnerabilities for Autonomous Driving Policy with Large Language Models

### v1

@@ V-B Evaluation of Diversity
To evaluate the capability of AED in discovering diverse vulnerability scenarios, we measure the number of distinct vulnerability types identified across multiple training iterations. For each traffic environment setting—defined by the traffic environment (Highway and Roundabout), the tested policy (VI, RVI, and D3QN), and the number of adversarial vehicle agents (2 and 3)—we conduct three independent training iterations with different random seeds. In each iteration, we collect all effective failure cases. In the first iteration, we record the most frequently occurring effective vulnerability type and assign it a count of 1. In the second and third iterations, we exclude types previously discovered and select the most frequently occurring type among the remaining types. Instead of simply incrementing the count by one, we add the relative frequency of the newly discovered type—its proportion among all effective failures in that iteration—to the cumulative count. In this way, each iteration contributes at most one type, and the final number reflects both the counts and the prominence of distinct vulnerability types discovered. For VDARS, each adversarial policy is trained using the same handcrafter reward function across all iterations. For AED, the adversarial policy in each iteration is trained using one of the three different reward functions, which are automatically generated by LLMs in parallel and targeted at different failure types (i.e., Left, Right, and Rear).
As shown in Fig. 3, AED consistently discovers a greater number of distinct vulnerability types compared to VDARS across all tested scenarios. In both the Highway and Roundabout environments, AED reaches or closely approaches the maximum of three types within three training iterations, regardless of the tested policy and the number of adversarial vehicle agents. In contrast, VDARS typically discovers only one or two types, even after multiple iterations. This discrepancy highlights the limited exploration capability of using a single handcrafted reward function in VDARS. On the other hand, AED benefits from multiple LLMs-driven reward functions that promote the exploration of different failure scenarios, enabling the adversarial agents to expose a broader range of policy weaknesses.
V-C Evaluation of Effectiveness
@@ VI RVI D3QN VI RVI D3QN
Highway Left VDARS $34.24\%_{(0.09)}$ $61.88\%_{(0.52)}$ $25.81\%_{(0.07)}$ $44.38\%_{(0.14)}$ $79.51\%_{(0.10)}$ $58.24\%_{(0.18)}$
@@
AED $\mathbf{100\%}_{(0.00)}$ $\mathbf{100\%}_{(0.00)}$ $\mathbf{90.00\%}_{(0.04)}$ $\mathbf{100\%}_{(0.00)}$ $\mathbf{100\%}_{(0.00)}$ $\mathbf{88.89\%}_{(0.05)}$
Right VDARS $39.56\%_{(0.11)}$ $11.18\%_{(0.09)}$ $0.00\%_{(0.00)}$ $26.92\%_{(0.08)}$ $5.13\%_{(0.01)}$ $0.00\%_{(0.00)}$
Eureka $90.00\%_{(0.00)}$ $83.89\%_{(0.28)}$ $90.56\%_{(0.39)}$ $\mathbf{92.78\%}_{(0.48)}$ $75.56\%_{(0.31)}$ $11.11\%_{(0.64)}$
AED $\mathbf{95.00\%}_{(0.04)}$ $\mathbf{87.77\%}_{(0.03)}$ $\mathbf{100\%}_{(0.00)}$ $\mathbf{90.00\%}_{(0.01)}$ $\mathbf{90.56\%}_{(0.10)}$ $\mathbf{12.22\%}_{(0.01)}$
Rear VDARS $2.28\%_{(0.03)}$ $1.49\%_{(0.01)}$ $1.94\%_{(0.01)}$ $1.45\%_{(0.01)}$ $0.85\%_{(0.01)}$ $0.00\%_{(0.00)}$

### v2

@@ INTRODUCTION
To evaluate AED’s ability to discover diverse vulnerability scenarios, we measure the number of distinct vulnerability types identified across multiple training iterations. We divide the three high-level collision categories into six fine-grained subtypes to enable more precise diversity assessment. Illustrative examples discovered in the Highway environment are shown in Fig. 4. A complete set of definitions and additional illustrative examples can be found in the Appendices. For each setting—defined by the environment (Highway and Roundabout), the tested policy (VI, RVI, and D3QN), and the number of adversarial vehicle agents (2 and 3)—we conduct six independent training iterations with different random seeds. In each iteration, we collect all effective failure cases. In the first iteration, we identify the most frequently occurring effective vulnerability type and assign it a count of 1. For each subsequent iteration, we exclude all previously discovered types and select the most frequent type among the remaining ones. Instead of simply incrementing the count by one, we add its proportion among all effective failures in that iteration to the cumulative count. This strategy ensures that each iteration contributes at most one new type, and the final diversity score reflects both the variety and prominence of distinct vulnerabilities uncovered. In VDARS, all iterations use the same handcrafted reward function. In contrast, AED trains the adversarial policy in each iteration with a different LLMs-generated reward function targeted at distinct failure types.
As shown in Fig. 5, AED consistently discovers more diverse vulnerability types than VDARS. Across all settings, AED reaches or approaches the maximum of six types within six iterations, while VDARS typically uncovers only two or three. This discrepancy highlights the limited exploration capability of a single handcrafted reward and AED’s strength in promoting broader vulnerability discovery through multiple LLMs-generated reward functions.
Evaluation of Effectiveness
We evaluate the effectiveness of the proposed preference-based enhancement module through both qualitative and quantitative analyses. These evaluations demonstrate its utility by: (1) assessing the learned reward model’s ability to distinguish goal-aligned vulnerabilities from unrelated failures, and (2) measuring its impact on the overall effective vulnerability rate when integrated into RL training.
Qualitative Evaluation
We first illustrate the reward model’s discriminative capability using a representative case from the Highway environment with three adversarial vehicle agents and a D3QN tested policy. When trained with a LLMs-generated reward function targeting Left Lane Change (Left) collisions, the resulting accidents include both targeted Left collisions, unintended Rear-End collisions, and other ineffective failures. After applying preference-based reward learning, the learned model is evaluated on different accident trajectories. As shown in Fig. 6(a), it assigns higher reward scores to Left collisions (orange) than to unrelated accidents (green), demonstrating clear separation. Similar reward distributions are observed across other environments and policies (Fig. 6(b)(c)(d)), further validating the model’s ability to distinguish goal-aligned vulnerabilities from irrelevant ones.
Figure 6: Probability density distribution of reward for accident trajectories across different environments and policy configurations. In all cases, the learned reward model assigns higher scores to goal-aligned failures (orange), effectively separating them from unrelated accidents (green).
Figure 7: Effective Vulnerability Rates (%) across different environments, tested policies, and numbers of adversarial vehicle agents. Each radar chart corresponds to one environment with one high-level accident category. Adv. is abbreviation for “adversarial”. AED consistently achieves the highest effective vulnerability rates.
Figure 8: Case study demonstrating the AED framework. (a) Task description provided to LLMs specifies the targeted Left Lane Change vulnerability, scenario details, and responsibility criteria. (b) The LLMs-generated reward function verifies the tested policy’s maneuver, ensures non-aggressive behavior from adversarial vehicles, and assigns responsibility accordingly. Differences from the expert-crafted reward are highlighted using a diff-style format. (c) The preference-based reward model differentiates targeted failure scenarios from unintended ones.
Quantitative Evaluation
To quantify the benefit of preference-based enhancement, we measure the effective vulnerability rate—defined as the percentage of simulation runs where the adversarial policy successfully triggers a specific failure type attributable to the tested policy. We compare the full AED framework against a baseline ablation, Parallel Eureka, which uses LLMs-generated reward functions without the final enhancement module. Parallel Eureka can be seen as a variant of VDARS, replacing a fixed handcrafted reward function with static LLMs-generated ones. In AED, if the learned reward model assigns a score below a predefined threshold (black dashed lines in Fig. 6), the corresponding LLMs-generated reward is suppressed to zero, thus filtering out off-target behaviors.
The effective vulnerability rate reported corresponds to the maximum among all fine-grained subtypes in each high-level collision category. Results for all subtypes are provided in the Appendices. As shown in Fig. 7, AED consistently outperforms or matches Parallel Eureka in nearly all settings. The improvement is especially notable where the initial LLMs-generated reward is suboptimal. For example, in the Highway environment with a D3QN-driven tested policy and three adversarial vehicles, AED increases the effective Left collision rate from $32.22\%$ to $88.89\%$. This substantial improvement demonstrates the impact of preference-based enhancement in guiding adversarial vehicles to consistently induce failure scenarios that the tested policy struggles with. Even in cases where Parallel Eureka already performs well, AED maintains or improves performance, indicating that the refinement module enhances reward fidelity without degrading strong initial signals.
Overall Performance
AED consistently outperforms VDARS in both diversity and effectiveness. It approaches a maximum of six distinct vulnerability types in most settings. It also consistently achieves significantly higher effective vulnerability rates (Fig. 7). These gains stem from AED’s design: its parallel accident generation enables broader exploration of failure modes, while the preference-based enhancement module refines noisy LLMs-generated rewards into precision training objects. Together, they make AED a more comprehensive framework for adversarial vulnerability discovery.
Interpretability and Discussion
To illustrate how AED transforms high-level vulnerability descriptions into effective adversarial policies, we present a case study in Fig. 8, targeting Left Lane Change collisions in the Highway environment with a D3QN-driven tested policy and three adversarial vehicles.
Task Specification and Reward Generation
We begin by prompting the LLMs with relevant environment context (e.g., road layout, lane width, vehicle length, and action set) and task description, including the targeted vulnerability type (“Left Lane Change Collisions”), expected behaviors (e.g., “the tested AV unexpectedly changes lanes to the left…”), and criteria for assigning responsibility. Based on this, the LLM generates $K$ ($K=6$ in our experiments) executable reward functions. Each candidate function is used to train an adversarial policy, and the best-performing one—measured by the proportion of Left Lane Change collisions in simulation runs—is selected.
Reward Function Analysis
The selected reward function is shown in Fig. 8(b). It consists of modular checks for: (1) whether the tested vehicle performed a left lane change before collision, (2) whether adversarial vehicles behaved non-aggressively (e.g, within steering angle and acceleration constraints), and (3) whether the collision can be attributed to the tested vehicle. The differences between the LLMs-generated reward and the expert-crafted version are visualized using a diff-style format. While both functions share a similar core structure—detecting collisions and evaluating adversarial passiveness—the expert reward relies on a simplified branching logic that only distinguishes between lane change and rear-end collisions. In contrast, the LLMs-generated reward introduces two key advantages. First, it explicitly encodes the targeted collision type (Left Lane Change) as part of the reward condition. Second, it examines both the current and previous actions of the AV to more accurately determine responsibility. These improvements ensure that rewarded collisions are not only aligned with the intended failure type, but also effectively attributed to decision-making flaws in the tested policy.
Preference-based Enhancement
Initial adversarial training with the LLMs-generated reward function yields a suboptimal distribution of accident types: Left Lane Change collisions $(36.67\%)$, Rear-End collisions $(1.00\%)$, and ineffective accidents $(62.33\%)$. As shown in Fig. 8(c), the preference-based model corrects this by assigning higher rewards to targeted Left Lane Change collisions and lower rewards to unintended or ineffective accidents. During RL training, the preference model is integrated with the original LLMs-generated reward function. If the preference score for a trajectory during training falls below a predefined threshold, the LLMs-generated reward is suppressed to zero. This filtering mechanism shifts the training focus towards high-quality, goal-aligned failures, resulting in a substantial increase in effective Left Lane Change collisions to $88.89\%$.
CONCLUSION
We propose AED, an LLMs-based framework that uses reinforcement learning to automatically discover effective and diverse vulnerabilities of autonomous driving policies. We use an LLM to automatically design reward functions without the need of human experts. To enhance the diversity of vulnerabilities, we use the LLM to consider a broad range of potential accidents and train adversarial policies using RL in parallel. We further improve the effectiveness of each vulnerability by training a reward model with preference-based learning to filter ineffective or other types of accidents. Experiments on various driving scenarios with both planning-based policies and learning-based policies show that AED discovers a diverse range of vulnerabilities and achieves higher attack success rates compared to existing methods, showing the potential of using LLMs to reduce the need for manual reward engineering and improve the comprehensiveness of autonomous driving safety evaluation.
Appendices
MAPPO Training
Observation and Action Space
In our experiments, all agents share the same observation and action space. Each agent receives an observation vector $o_{i}=[k_{0}^{i},k_{1}^{i},k_{2}^{i},k_{3}^{i},k_{4}^{i}]$, where each $k_{j}^{i}$ represents the state of one of the five nearest neighboring vehicles. Each $k$ consists of five features $k=[p,x,y,v_{x},v_{y}]$ as detailed in Table. 2. Each agent selects from a discrete action space consisting of five maneuver options, as shown in Table. 2.
Feature Description Range
$p$ Binary indicator of vehicle presence $\{0,1\}$
$x$ Relative longitudinal position $150,150]\mathrm{m}$
$y$ Relative lateral position $16,16]\mathrm{m}$
$v_{x}$ Relative velocity along x-axis $60,60]\mathrm{m/s}$
$v_{y}$ Relative velocity along y-axis $60,60]\mathrm{m/s}$
Table 1: Observation space in MAPPO training.
Index Discription
0 Change to the left lane
1 Maintain current lane and speed
2 Change to the right lane
3 Accelerate
4 Decelerate
Table 2: Discrete action space available to agents.
Environment Setting
To account for diverse and realistic traffic scenarios, we conduct experiments in two representative simulation environments: “Highway” and “Roundabout’ as demonstrated in Fig. 9. Highway scenarios involve high-speed interactions, frequent lane changes, rear-end collisions and overtake collisions, which are typical in structured multi-lane traffic. Roundabout scenarios involve frequent turning, right-of-way decisions and merging behaviors. The key configuration parameters are summarized in Table. 3, which define the road structure, background vehicle behavior, and task-specific constraints.
Figure 9: Demonstration of “Highway” and “Roundabout”.
Parameter Value
Number of Lanes 4
Number of Background Vehicles 5
Background Vehicle Control Policy Value Iteration
[…]
Max Task Length 20
Table 3: Environment Configurations used in MAPPO training.
Training Hyperparameters
Table 4 summarizes the hyperparameters used in our MAPPO training for adversarial vehicle agents. These configurations are kept consistent across all environments.
Hyperparameter Value
Network Architecture MLP+RNN
Hidden Layer Size 64
Initial Learning Rate 5e-4
Discount Factor 0.99
GAE Parameter 0.95
Gradient clipping 0.5
Value Loss Coefficient 0.5
Entropy Coefficient 0.01
PPO Clipping Parameter 0.2
PPO Epochs 15
Training Steps 10M
Rollout Threads 100
Max Episode Length 40
Table 4: Hyperparameters used in MAPPO training.
Figure 10: Examples of distinct vulnerability types discovered by AED. Adv. is abbreviation for “adversarial”.
Figure 11: Example of T-Bone collision.
Illustrations of Discovered Vulnerability Scenarios
To better reflect real-world traffic incident patterns, we further decompose the three high-level categories—Left Lane Change, Right Lane Change, and Rear End collisions—into six fine-grained subtypes. These subtypes enable a more detailed assessment of vulnerability diversity:
  • Left-Sparse: The tested vehicle performs a left lane change without properly estimating the distance of an approaching adversarial vehicle in the adjacent lane, resulting in a sideswipe collision in sparse traffic.
  • Right-Sparse: The tested policy executes a right lane change while misjudging the lateral distance to a nearby adjacent vehicle, causing under sparse traffic conditions.
  • Rear End: The tested vehicle fails to maintain a safe following distance or reacts too slowly, resulting in a rear-end collision with an adversarial vehicle in the same lane.
  • Overtake: The tested policy attempts to overtake a slower adversarial vehicle ahead, but misjudges the timing and crashes during the maneuver, typically due to its failure to account for adversarial vehicles blocking the return lane.
  • Left-Dense: When facing congestion caused by an adversarial vehicle ahead, the tested policy attempts a left lane change but collides with another adversarial vehicle in the adjacent lane.
  • Right-Dense: Similarly, due to congestion caused by a leading adversarial vehicle, the tested policy attempts a right lane change and causes a sideswipe collision.
Fig. 10 shows example trajectories for each of the above vulnerability subtypes in the Highway Environment. In the Roundabout environment, overtaking accidents are really rare in real-world scenarios and are therefore excluded from our analysis. Instead, we identify a distinct type of vulnerability unique to this setting as shown in Fig. 11:
  • T-Bone: The tested vehicle misjudges the right-of-way when entering the roundabout, and fails to yield to an adversarial vehicle already circulating within it. This results in a perpendicular collision where the front of the tested vehicle crashes into the side of another.
@@ VI RVI D3QN VI RVI D3QN
Highway Left-Sparse VDARS $34.24\%_{(0.09)}$ $61.88\%_{(0.52)}$ $25.81\%_{(0.07)}$ $44.38\%_{(0.14)}$ $79.51\%_{(0.10)}$ $58.24\%_{(0.18)}$
@@
[…]
Right-Sparse VDARS $39.56\%_{(0.11)}$ $11.18\%_{(0.09)}$ $0.00\%_{(0.00)}$ $26.92\%_{(0.08)}$ $5.13\%_{(0.01)}$ $0.00\%_{(0.00)}$
[…]
Rear VDARS $2.28\%_{(0.03)}$ $1.49\%_{(0.01)}$ $1.94\%_{(0.01)}$ $1.45\%_{(0.01)}$ $85\%_{(0.01)} 0.0$$00\%_{(0.00)}$